OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 3x2 + 5x + 1 = 0

pooja195 (pooja195):

list your abc values consider: \[ax^2+bx+c=0\]

OpenStudy (anonymous):

\[\frac{ -5\pm \sqrt{5^{2}-4(3)(1)} }{ 2(3) }\]

OpenStudy (abdullahm):

Correct, now just simplify that.

OpenStudy (anonymous):

\[\frac{ -5\pm \sqrt{13} }{ 6 }\]

OpenStudy (nurali):

Correct

OpenStudy (abdullahm):

Good job :)

OpenStudy (nurali):

x = (-5 +/- 3.6) / 6 x = (-5 - 3.6)/6 or x = (-5 + 3.6)/6 x = -8.6/6 or x = -1.4/6 x = -1.43 or x = -0.23

OpenStudy (anonymous):

ur welcome